OpenStudy (anonymous):

Solve this system using substitution. Write your answer as an ordered pair. x = -3 3x – 4y = -37

Mehek (mehek14):

substitute \(x=-3\) into the second equation

Mehek (mehek14):

\(3(-3)-4y = -37\)

Mehek (mehek14):

what is \(3*-3?\)

Mehek (mehek14):

@Spencerstrickling18 ?

OpenStudy (anonymous):

3*-3 is -9

OpenStudy (anonymous):

-9-4y = -37 -28 = -4y 7 = y ordered pair is in form (x,y)